OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

@tuugongomdal when two lines are parallel, their slopes are equal in order to find the eqn passing through points (x1,y1) , use slope point form which is defined as y-y1 = m (x-x1)

OpenStudy (anonymous):

what do i do?

OpenStudy (anonymous):

do u know what do u want to determine in the problem? here u r not gonna get direct solutions. U have to work with us. BTW welcome to this amazing world of openstudy

OpenStudy (anonymous):

thanks and are you asking for the possible answers?

OpenStudy (anonymous):

no i m not asking for the possible answers but i want to know how did u approach the problem so that u can be helped in a better way

OpenStudy (anonymous):

i dont understand what you mean by approach the problem. i have no idea how to do it at all

OpenStudy (anonymous):

seems u didnot attempt the ques

OpenStudy (anonymous):

i cant attempt something i have not a clue to do, thats why im here so i can be taught

OpenStudy (anonymous):

u must have learn some concepts, thats y u r doing problems

OpenStudy (anonymous):

lets do this ques hope i can help u

OpenStudy (anonymous):

in the problem , equation of the line is given . U need to determine the equation that is parallel to the line and passing through some point equation of the line y = 2x + 2 is in the form of slope intercept form which is defined as y = mx+c where m is slope of the line can u determine the slope of the line?

OpenStudy (anonymous):

i dont remember how

OpenStudy (anonymous):

u need to compare the equation of the given line with slope intercept form

OpenStudy (anonymous):

nevermind your not helping very well

OpenStudy (anonymous):

it seems so i think other ppl will help u in the better way

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

one common form of a straight line is \[y=\color{brown}mx+\color{teal}b\] where \(\color{brown}m\) is the slope or gradient and \(\color{teal}b\) is \(y\)-intercept can you tell me the slope of your line? \[y = \color{brown}2x + \color{teal}2\]

OpenStudy (unklerhaukus):

line that are parallel have the same \(\color{brown}{\text{slope}}\)

OpenStudy (unklerhaukus):

once you have the slope \(\color{brown} m \), plug in the co-ordinates \((\color{orange}5, \color{green}3)\) \[\color{green}y=\color{brown}m\color{orange}x+\color{teal}b\] and solve the equation for \(\color{teal}b\) . Your final solution wil be \[y=\color{brown}mx+\color{teal}b\]